OpenStudy (anonymous):

Calculate the antiderivative of dx/(x^2-9)

OpenStudy (ash2326):

We have \[\int \frac{dx}{x^2-9}\]\ this can be written as \[ \int \frac{dx}{(x+3)(x-3)}\] Let's multiply and divide the numerator by 6 \[ \int \frac{6dx}{6(x+3)(x-3)}\] Now 6 can be written as \[ 6=(x+3)-(x-3)\] Now the numerator can be replaced by (x+3)-(x-3) we get \[ \int \frac{(x+3)-(x-3)}{6(x+3)(x-3)}dx\] We get now \[ \int \frac{dx}{6(x-3)}- \frac{dx}{6(x+3)}\] Now performing integration \[ \frac{1}{6}\ln(x-3)-\frac{1}{6}\ln(x+3)+c\] Log terms can be combined , we get \[\frac{1}{6}\ln{\frac{x-3}{x+3}}+c\]

OpenStudy (anonymous):

...Very clever, Ash. I was going to suggest a trigonometric substitution of sin(x), but looks like you gots it, yo.

OpenStudy (ash2326):

Thanks Badreferences:D

OpenStudy (anonymous):

I'm GOING to have to remember that trick.

OpenStudy (ash2326):

Caro did you understand it?

OpenStudy (anonymous):

If the asker is curious, if you use the identity cos(x)^2=1-sin(x)^2, and you use x=nsin(x), you can solve this a la trigonometric substitution.

OpenStudy (anonymous):

Yup that was a perfect/clear explanation. Thans a lot!!